Find the equation of the line which is perpendicular to 4𝑦=5π‘₯βˆ’8 and passing through (2,3).


1
Expert's answer
2022-02-07T17:47:56-0500
"4𝑦=5𝑥−8=>y=\dfrac{5}{4}x-2"

"slope_1=m_1=\dfrac{5}{4}"

If two lines with slopes "m_1" and "m_2" are perpendicular then "m_1m_2=-1."

Then "\dfrac{5}{4}m_2=-1."


"slope_2=m_2=-\dfrac{4}{5}"

The equation of the second line is


"y=-\dfrac{4}{5}x+b"

This line passes through the point "(2, 3)"


"3=-\dfrac{4}{5}(2)+b=>b=y=\dfrac{23}{5}x"

The equation of the line which is perpendicular to "4\ud835\udc66=5\ud835\udc65\u22128" and passing through "(2,3)" is


"y=-\dfrac{4}{5}x+\dfrac{23}{5}"

Or


"5y=-4x+23"